OpenStudy (anonymous):

How do i change x+3y=6 to y=mx+b?

OpenStudy (anonymous):

You have to arrange the equation. So its y=mx+b. It would turn into 3y+x=6. Then you have to get Y by itself; so you divide by 3 on both sides. It would then turn into y=3x+2

OpenStudy (anonymous):

x+3y=6 (subtract x on both sides) 3y=-x+6 (divide by 3 on both sides) y=(1/3)x+2

OpenStudy (anonymous):

Thank you